o1-The angular position of a angular speed of the point at t = 3s (A)6 rad s(B) 10 rads (C)14r point on a wheel is described by θ (t) = 2t + 2t3(rad). Find t ad s-t (D) 18 rad s1 (E) 22 reds A wheel starts from rest and rotates with const ant angular acceleration to reach an ingular spec
Answer #1

angular speed = rate of change of angular position

w = d\theta / dt

= d(2t + 2t^2) / dt

= 2 + 4t

so w(t) = 2 + 4t

w(3) = 2+ 4*3 = 14 rad/s

so option C) is the correct answer